# Efficient Input-Constrained Impulsive Optimal Control of Linear Systems with Application to Spacecraft Relative Motion

### v1

@@ V-C VISORS Transfer Demonstration
The VISORS mission [22] is a two-nanosatellite mission that aims to perform telescopic measurements of the solar corona. Part of the mission operations involves performing a long duration, impulse constrained transfer from a 200 meter radius standby orbit to a 40 meter science orbit. The challenge of this transfer is confounded by the limitations of the propulsion systems aboard the nanosatellites, which can provide a maximum impulse of just 2 millimeters per second with at least 45 seconds of cool down time between each impulse. The current technique used to realize this transfer involves applying closed-form solvers to a sequence of manually specified waypoints that gradually tightens the relative orbit. The main issues with this approach are that it is inflexible to incorporating dynamical perturbations, it does not provide optimal solutions, and it is not adaptable to differing initial and final conditions.
@@
To demonstrate the strength of the proposed algorithm for this difficult problem, the standby to science scenario from [22] is implemented. Initial and final conditions for this transfer are given in Table II. The transfer duration is 10 orbits and 500 equally spaced time windows are enforced with a buffer of 45 seconds and a maximum impulse expenditure of $2$ millimeters per second per window. The $L_{1}$ norm is applied such that $f_{k}(\cdot)=\|\cdot\|_{1}\ \forall k$. Each time window is discretized over 100 time points. Finally, the orbital mean motion is $n=.00113$.
TABLE II: Initial Conditions for VISORS Reconfiguration (m)
$a\delta a$ $a\delta\lambda$ $a\delta e_{x}$ $a\delta e_{y}$ $a\delta i_{x}$ $a\delta i_{y}$
$\delta\textit{{\oe }}_{0}$ 0 0 0 200 0 200
$\delta\textit{{\oe }}_{f}$ -2.62 45.21 -34.51 4.78 -18.72 2.72
(a) Relative Position
(b) $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$
Figure 3: VISORS Reconfiguration Trajectory
(a) Throttle Profile
(b) Dual Norm Profile
Figure 4: VISORS Reconfiguration Input Profile
(a) $\delta a-\delta\lambda$
(b) $\delta\bm{e},\delta\bm{i}$
(c) RN Separation
Figure 5: VISORS Reconfiguration Relative Orbital Elements
The resulting trajectory is displayed in Figure 3, the input profile in Figure 4, and the relative orbital element perspective in Figure 5. In Figure 3b, the trajectory of $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$, equivalent to the primer vector, is plotted along with the unit cube, which corresponds to the unit sublevel set of the dual norm. Times when the primer vector leaves this region correspond to possible impulse events. The dual norm of the primer vector is shown in Figure 4b, along with time windows and impulse times, which occur when the dual norm signal is maximized over the time window. The relative orbital elements reveal an important feature of interest of the computed maneuver sequence, namely that RN separation, plotted in Figure 5c and typically used to evaluate the safe spacecraft separation [28], is maintained above 18 meters for the entire transfer duration, despite the fact that no state constraints are applied in the solution procedure.
The maneuver sequence under the proposed algorithm requires 0.3095 meters per second of total $\Delta V$ expenditure. In contrast, [22] reports 0.4934 meters per second of $\Delta V$ expenditure for the same transfer. Given that this transfer is performed 20 times over the mission span, this amounts to a total $\Delta V$ saving of 3.68 meters per second, which is $\sim$25% of the total $\Delta V$ budget of the CubeSat. The proposed algorithm is additionally capable of incorporating $J_{2}$ and drag perturbations through the linear model formulations in [25], does not require manual tuning of waypoints along the trajectory, and is flexible to changing initial and final conditions. The total solver clock time is 0.2675 seconds on a modern computer, which is reasonably achievable in an online manner under the lower computational resources of the VISORS CubeSats.
@@ VI Conclusion
This paper demonstrated an extension to a class of moment problems that included constraints on the integral of the norm of the input signal over time windows of interest. It was shown that the resulting optimal solution remains impulsive and that semi-infinite programming can be applied to global optimality and global convergence. This algorithmic framework was applied to challenging spacecraft relative motion control problems, where it was used to derive maneuver sequences that outperformed established techniques in terms of optimality.
Though this paper has focused on applications to linear time-varying systems, the authors believe that the theoretical framework outlined has significant promise for extending indirect trajectory optimization techniques for impulsively controlled nonlinear systems. Future work could also include the incorporation of state constraints and extensions to impulsive covariance control.

### v2

@@ V-C VISORS Transfer Demonstration and Comparison
The VISORS mission [11] is a two-nanosatellite formation-flying mission that aims to perform telescopic measurements of the solar corona. Part of the mission operations involves performing a long duration, impulse constrained transfer from a 200 meter radius standby orbit to a 40 meter science orbit, i.e., $a\delta\textit{{\oe }}_{0}=[0,0,0,200,0,200]^{\intercal}$ [m] to $a\delta\textit{{\oe }}_{f}=[-2.62,45.21,-34.51,4.78,-18.72,2.72]^{\intercal}$ [m]. The challenge of this transfer is compounded by the limitations of the propulsion systems aboard the nanosatellites, which can provide a maximum impulse of just 2 millimeters per second with at least 45 seconds of cool down time between each impulse.
@@
To demonstrate the strength of the proposed algorithm for this difficult problem, the standby to science scenario from [11] is demonstrated with a transfer duration of 10 orbits and 500 equally spaced time windows are enforced with a buffer of 45 seconds and a maximum impulse expenditure of $2$ millimeters per second per window. The $l_{1}$ norm is applied, $f_{k}(\cdot)=\|\cdot\|_{1}\ \forall k$, to be consistent with the fuel metric in [11]. Each time window is discretized over 100 time points.
Figure 3: VISORS Transfer Trajectory.
Figure 4: VISORS Transfer Relative Orbital Elements
The computed VISORS transfer trajectory is displayed in Figure 3, including the RTN trajectory, a subset of the impulse profile, and a subset of the dual norm trajectory. Additionally, the relative orbital element trajectory is plotted in Figure 4, along with the RN plane separation, which validates safe spacecraft separation over the transfer duration [8].
Finally, this scheme is embedded in a high-fidelity simulation, with the inclusion of execution errors, navigation uncertainty, and dynamical effects that are consistent with [11]. The control dynamical model includes drag and $J_{2}$ effects, as derived in [15], and the transfer is replanned every orbit for the first 9 orbits and every quarter orbit during the final orbit. Table I summarizes the performance over 250 Monte Carlo trials. Importantly, the proposed approach results in an average of $33.27$ [cm/s] of $\Delta V$ expenditure for the transfer, which substantially improves on the $49.34$ [cm/s] metric reported in [11]. Given that this transfer is performed 20 times over the mission span, this amounts to an estimated total $\Delta V$ saving of 3.19 meters per second, which is $\sim$25% of the total $\Delta V$ budget of the VISORS CubeSats.
TABLE I: VISORS Transfer Performance over 250 Monte Carlo Trials with $1\sigma$ Bounds
Total $\Delta V$ Proposed Solver Time Direct Solver Time
$33.27\pm 0.72$ [cm/s] $0.1912\pm 0.0480$ [s] $0.4234\pm 0.0263$ [s]
@@ VI Conclusion
This paper demonstrated an extension to a class of moment problems that includes constraints on the integral of the norm of the input signal over time windows of interest. It was shown that the resulting optimal solution remains impulsive and that semi-infinite programming can be applied to obtain optimal solutions with guaranteed convergence. This algorithmic framework was applied to challenging spacecraft relative motion control problems, where it was used to compute maneuver sequences that outperformed established techniques in terms of optimality and runtime.